(D^4 − 18D^3 + 119D^2 − 342D + 360)y = 0

1
Expert's answer
2022-05-15T11:30:39-0400

Solution

For given homogeneous equation the characteristic equation is

λ4 – 18 λ3 + 119 λ2– 342 λ + 360 = 0

λ4 – 3 λ3 – 15 λ3 + 45 λ2 + 74 λ2 – 222 λ – 120 λ + 360 = 0

(λ – 3)( λ3 – 15 λ2 + 74 λ – 120 ) = 0

(λ – 3)( λ3 – 4 λ2 – 11 λ2 + 44 λ + 30 λ – 120 ) = 0

(λ – 3) (λ – 4) ( λ2 – 11 λ + 30 ) = 0

(λ – 3) (λ – 4) (λ – 5) (λ – 6) = 0

So solutions of this equation are λ1 = 3, λ2 = 4, λ3 = 5, λ4 = 6.

Therefore the general solution of the given differential equation is

y(x) = Ae3x + Be4x + Ce5x + De6x , where A, B, C, D are arbitrary constants. 

Answer

y(x) = Ae3x + Be4x + Ce5x + De6x